John invested 3 times as much money at 12% interest as he did at 8%. If he earned a total of $264 in interest for the year, how much did he invest at each rate?
Answer by addingup(3677)

Invested at 8%: x
Invested at 12%: 3x
0.08x+0.12(3x) = 264
0.08x+0.36x = 264
0.44x = 264
x = 600
3x = 1800
He invested 600 at 8% and 600*3 = 1800 at 12%
--------------------------------
Check:
600*.08 = 48
1800*0.12 = 216
48+216= 264 correct